# We Have a Package for You! A Comprehensive Analysis of Package Hallucinations by Code Generating LLMs

### v2

@@ 5.4 Characteristics of Hallucinations (RQ4)
Occurrence of the Same Package Hallucination Across Different Models. To analyze the possibility of the same hallucinated packages being generated across different models, we measured how many models generated the same package name given a confirmed package hallucination. Figure 8 shows a clear pattern where a vast majority (81%) of distinctly generated package names were generated by only one model. In other words, specific package names were usually unique to a single model, where only the most common packages were generated by more than one model. The two populations diverge as the number of models increases, with the number of hallucinated packages decreasing nearly exponentially and the distribution of valid packages becoming more uniform. The finding that valid packages are less dependent on the specific model used for generation is attributable to their frequent appearance in training data and applicability to universal coding problems, leading to their widespread use in a broad range of prompts.
Combining the insights gained during the previously discussed persistence analysis (Figure 5) leads to a key observation. As previously shown, hallucinations are often persistent (58% are repeated within 10 iterations) within the same model but are not often repeated between models, as 81% of hallucinated packages are generated by only one model. This further reinforces the evidence that while hallucinations are a common phenomenon across various models, the exact nature of these hallucinations is generally model-specific. This behavior is particularly surprising given that our testing includes multiple models from the same family (i.e. 3 GPT models, 4 CodeLlama models, and 3 DeepSeek models). These models presumably use the same training data for each version, yet each model generates unique hallucinations not found in other models.
@@
Semantic Similarity Between Hallucinated and popular packages. In order to analyze the semantic similarity between hallucinated and popular real/valid packages, we measured the average Levenshtein distance of a package to its nearest neighbor (i.e., the closest valid package). Levenshtein distance is a measure of how many insertions, deletions, and substitutions are required for two strings to match [36]. If the distribution of Levenshtein distances is skewed heavily right, with a peak at or near 0, this would indicate that most hallucinations are very similar to valid package names. In that case, attackers could infer a hallucination target based on more traditional package confusion methods (e.g., typosquatting) rather than analyzing a large volume of model output over time to detect persistent hallucinations that could be used as vessels for malicious code. A higher distance reflects that package hallucinations are more random in nature and difficult to predict, rather than the result of minor grammatical errors.
@@
The results of the Levenshtein distance test, as seen in Figure 9, suggests that most package hallucinations are not simple off-by-one errors. An off-by-one error in our case is defined as a difference between the hallucinated package and its closest match of 1-2 numbers, letters, or punctuation marks, with only 13.4% (10,263 of 76,489) belonging to this category. Another 37.9% (29,025 of 76,489) of packages registered a score between 3-5, which would indicate two words with a common root word or concept that still differ significantly. Notably, 48.6% (37,207 of 76,489) of hallucinations scored 6 or greater, with 20.2% (15,457 of 76,489) of those scoring 10 or higher, indicating two strings that are very different and likely do not share any common theme.
@@
The presence of such a large proportion of high Levenshtein values suggests that the majority of hallucinations are not merely trivial typographical errors but are instead substantively different from existing package names. The observed results provide further evidence that the root cause of hallucinations is likely to be more complex than minor string manipulation, pointing to deeper issues in the model’s generative processes that govern the creation of package names. The long right tail of the distribution in Figure 9 indicates a wide variety of hallucinations spread across a broad range, revealing a diversity in types of errors and reinforcing that the generation of hallucinations is a complex issue not limited to simple character substitutions, additions, and deletions.
@@
Effect of Deleted Packages. To determine whether packages that existed before the pre-training data cut-off date (i.e., final day of data included in the model’s training set) but were subsequently removed, contribute significantly to package hallucinations, we conducted an analysis using package download counts obtained via Google BigQuery [16]. We searched PyPI download counts from 2022 and earlier to compile a list of packages that existed before 2023 but are no longer available on PyPI. This list was then compared against the master list of packages (obtained from PyPI as of January 10, 2024) of hallucinated packages across all models. We detected 12,871 packages that were available between 2020 and 2022 that have since been removed from PyPI. Out of these deleted packages, only 133 (0.17%) were generated during our analysis, indicating that deleted packages are a negligible source of package hallucinations. This finding contradicts our hypothesis, as we expected a sizable percentage of hallucinated packages due to the presence of deleted packages in the training data.
@@
Effect of Language Confusion. Another behavior observed during the main experiment is the tendency to confuse programming languages while generating package output (i.e. the model is asked to provide Python packages, but instead provides JavaScript packages). To validate the potential existence of this behavior in code generating LLMs, we obtained master lists of packages (using libraries.io [72]) from the nine most popular open-source repositories and compared our list of hallucinated package names generated during Python testing to the respective master lists of valid packages from other programming languages. Any intersection between the two lists indicates a cross-language hallucination. Overall, only JavaScript is a significant source of cross-language hallucinations, as 8.7% (6,705/76,489) of hallucinated Python packages are valid JavaScript packages. All other languages contributed negligible hallucinations, combining for only 0.8% (663 of 76,489) across eight other open-source repositories, including R, Rust, Ruby, PHP, Swift and .NET (see Table 3 for complete results).
@@
RQ4 Summary: Most unique hallucinated package names were generated by a single model and thus only appeared in 1 out of 16 tests. Most hallucinated package names were not semantically similar to a valid Python package as measured by Levenshtein distance. Deleted packages were a negligible source of package hallucinations while JavaScript was the only significant source of cross-language hallucinations.
@@ 6 Mitigation
Motivated by the finding of systemic package hallucination in all the tested models, we investigate techniques to mitigate the occurrence of such hallucinations (thus addressing RQ5).
@@ 6.1 Mitigation Strategies
A seemingly obvious mitigation strategy when detecting package hallucinations might involve comparing a master list of valid packages to the model’s code output. It is imperative to understand why this approach is merely a superficial solution rather than addressing the root cause. The fundamental vulnerability inherent in code generation by LLMs regarding package hallucinations is not solely that they produce non-existent packages, but that they generate package names that were not present in their training data. Our study emphasizes the detection of hallucinated packages as indicators of an underlying vulnerability. The genuine threat posed by package hallucinations emerges when such a hallucinated package is officially published and contains malicious code. The mere act of publishing does not render a previously hallucinated package benign; on the contrary, it becomes an active threat seeking to exploit a security vulnerability. Consequently, simply cross-referencing generated packages with a master list would erroneously validate a malicious package after publication. Dedicated mitigation techniques are essential to prevent the generation of hallucinated packages altogether, thereby addressing the root cause of this latent vulnerability.
Implementing mitigation strategies specific to code generation LLMs is an unexplored topic at the time of writing, therefore we rely on general hallucination mitigation strategies proposed for standard NLP tasks that can be applied to code generation. These strategies can generally be grouped into two broad categories: prompt engineering and model development [73]. Prompt engineering includes methods such as Retrieval Augmented Generation (RAG), self-refinement, and prompt tuning. RAG approaches involve enriching the original prompt with additional information gathered from an external source, such as the web or a pre-determined database [37]. This augmentation can occur at any stage—before, during, or after response generation—and can be iterative, improving over multiple cycles until the response is verified as accurate. Self-refinement strategies, on the other hand, utilize the model itself to detect and refine potential hallucinations.
The second main mitigation strategy involves improving the underlying LLM model itself through improved decoding strategies or supervised fine-tuning. Supervised fine-tuning alters model parameters to improve performance on tasks prone to hallucinations, utilizing a labeled dataset for more precise training. Decoding strategies are also considered a viable mitigation strategy from the literature, but based on our findings from RQ2 we know that altering decoding parameters, such as top-$k$, top-$p$, and min-$p$, do not result in a decreased hallucination rate. We evaluate each of the three remaining categories to determine their applicability to code generation and our specific use case of reducing hallucinations, ultimately developing a viable mitigation strategy.
@@ 6.2 Mitigation Implementation and Results
Retrieval Augmented Generation (RAG). We utilized a method of before generation RAG to supplement the prompt with valid package names to assist the model in generating a response. We made an additional dataset to serve as the supplementary information by taking the top 20,000 most popular packages from PyPI and prompting LlaMA-2 to generate a list of 5 questions that each package could help answer given the description. After removing duplicate responses, this resulted in 65,000 statements in the form “Package [x] could answer questions about [y]”. These 65,000 statements were stored in a vector database, which enables efficient retrieval of semantically similar statements. When a model was asked to recommend packages given a code generation prompt or Stack Overflow question, the vector database is first queried from within the code to return the top 5 most semantically similar statements. These statements are appended to the prompt to give the model additional information containing established valid packages to generate a non-hallucinated response.
Self-Refinement. Drawing on insights from our findings in RQ3 (see Section 5.3), which revealed that LLMs often exhibit proficiency in identifying their own package hallucinations, we implemented a self-refinement method. Following the generation of package names, the model is queried regarding the validity of these packages. If the model indicates that the packages are invalid, the response is regenerated with a specific instruction to not use the invalid package. This regeneration process is allowed to iterate up to five times, acknowledging that many package hallucinations are persistent, as demonstrated in RQ3, and may be repeatedly generated. It is possible during this test that a valid package is misclassified as a hallucination by the model, although because this is an iterative process the success rate should outweigh the false positive rate over the course of testing, resulting in a net decrease in hallucination rate.
Fine-tuning. For our next method, we fine-tuned the models using the code/package list (Heuristic 1) and prompt/package list (Heuristic 2) pairs that were generated during our initial experiments (Section 5.1). All hallucinations were filtered out and the models were re-trained using the remaining valid responses (560,000 samples). As fine-tuning affects the underlying model weights, we also need to make sure the fine-tuned model retains the ability to produce functional and effective code, which we will test by comparing the code quality from the original models and the fine-tuned models using the well-known HumanEval benchmark [8]. The HumalEval benchmark is a set of pre-defined prompts and test cases and the final score is a percentage of problems for which the model generated code is both syntactically correct and passes the test cases, reflecting the model’s ability to produce functional and accurate code.
@@
Results. Overall, all the mitigation strategies we implemented resulted in a reduced rate of package hallucination, with RAG and Supervised Fine-Tuning proving particularly effective (see Table 4). Fine-tuning proved to significantly improve results, especially for the DeepSeek model, where hallucinations were reduced by 83%, achieving a total rate of just 2.66%, which is a lower rate than any of the ChatGPT models (observed during RQ1). Self-refinement feedback was also much more effective for the DeepSeek model (19% reduction) compared to the CodeLlama model (3% reduction). This aligns with our results in RQ3, where the DeepSeek model was proficient at detecting hallucinations while the CodeLlama model had a strong bias towards labeling packages as valid, that limited its ability to reliably detect errors. The ensemble method of combining all mitigation strategies further improved results, reducing hallucination rates by 85% and 64% from their baseline levels for DeepSeek and CodeLlama, respectively.
The code quality of the fine-tuned models did decrease significantly, -26.1% and -3.1% for DeepSeek and CodeLlama respectively, in exchange for substantial improvements in package hallucination rate. While the code quality was negatively impacted, the fine-tuned scores are still at comparable levels to other high performing models such as Mistral 7B (26.1%), Llama 65B (23.1%), and Llama 2 7B (12.8%).
@@ 7 Discussion and Conclusion
One limitation of our study is the emergence of more advanced models since our evaluations. These newer models may offer improved performance and different hallucination characteristics, which were not captured in our study. The study also includes fewer commercial models due to funding constraints, meaning the findings may not fully represent the performance and hallucination tendencies of the latest commercial LLMs.
In terms of future work, the precise underlying causes of package hallucinations is still an open question. This includes exploring the architecture and components of LLMs that may contribute to these errors, examining the adequacy of tokenizers, and assessing how training data composition and preprocessing impact hallucination tendencies. Identifying and mitigating these underlying issues could lead to more robust and reliable code generation models. Future work could also focus on developing and testing more sophisticated mitigation strategies tailored specifically for code generation tasks. This could involve advanced techniques in prompt engineering, leveraging complex knowledge graphs, refining loss functions, and exploring new fine-tuning methods. Integrating real-time feedback mechanisms to dynamically adjust model output could further reduce hallucination rates. Understanding how package hallucinations are systemic and persistent at the token level remains crucial.
In conclusion, we systematically studied package hallucinations in code generation LLMs, including both commercial and open-source models. Our comprehensive analysis revealed that 19.7% of generated packages are fictitious, posing a critical threat to software security through package confusion attacks. We identified key behavioral patterns and characterized hallucinated packages, proposing effective mitigation strategies. Our findings underscore the importance of addressing package hallucinations to enhance the reliability and security of AI-assisted software development.
@@
We have disclosed our research to model providers including OpenAI, Meta, DeepSeek, and Mistral AI. As of this writing we have received no response or feedback. Our research highlights the feasibility of a new attack vector that can be used to carry out package confusion attacks by exploiting the package hallucinations generated by code-generating LLMs. Hallucinations are a well-studied limitation of generative AI models, including code-generating LLMs, which even current state-of-the-art techniques cannot completely eliminate. Our goal in this research is to get a better understanding of this phenomenon of package hallucinations in code-generating LLMs. We hope that these findings can help secure future models so that users are better protected against package confusion threats enabled due to package hallucinations.
@@
We are committed to the open science policy and will make all source code, datasets, and generated code publicly available. The only exception is we will not release our master list of hallucinated package names to the public. Such information could be misused by malicious actors to execute a package hallucination attack. The master list will be shared responsibly by request to verified researchers.

### v3

@@ 5.4 Characteristics of Hallucinations (RQ4)
Occurrence of the Same Package Hallucination Across Different Models. To analyze the possibility that the same hallucinated packages are generated across different models, we measured how many models generated the same package name given a confirmed package hallucination. Figure 8 shows a clear pattern in which a large majority (81%) of distinctly generated package names were generated by only one model. In other words, the specific package names were usually unique to a single model, where only the most common packages were generated by more than one model. The two populations (valid and hallucinated packages) diverge as the number of models increases, with the number of hallucinated packages decreasing nearly exponentially and the distribution of valid packages becoming more uniform. The finding that valid packages are less dependent on the specific model used for generation is attributable to their frequent appearance in training data and applicability to universal coding problems, leading to their widespread use in a broad range of prompts.
Combining the insights gained during the persistence analysis (Figure 5) leads to a key observation. As previously shown, hallucinations are often persistent (58% are repeated within 10 iterations) within the same model but are not often repeated between models, as 81% of hallucinated packages are generated by only one model. This further reinforces the evidence that while hallucinations are a common phenomenon across various models, the exact nature of these hallucinations is generally model specific. This behavior is particularly surprising given that our testing includes multiple models from the same family (i.e. 3 GPT models, 4 CodeLlama models, and 3 DeepSeek models). These models presumably use the same training data for each version, yet each model generates unique hallucinations that are not found in other models.
@@
Semantic Similarity Between Hallucinated and Popular Valid Packages. In order to analyze the semantic similarity between hallucinated and popular real/valid packages, we measured the average Levenshtein distance of a package to its nearest neighbor (i.e., the closest valid package). Levenshtein distance is a measure of how many insertions, deletions, and substitutions are required for two strings to match [31]. If the distribution of Levenshtein distances is skewed heavily right, with a peak at or near 0, this would indicate that most hallucinations are very similar to valid package names. In that case, attackers could infer a hallucination target based on more traditional package confusion methods (e.g., typosquatting) rather than analyzing a large volume of model output over time to detect persistent hallucinations that could be used as vessels for malicious code. A higher distance reflects that package hallucinations are more random in nature and difficult to predict, rather than the result of minor grammatical errors.
@@
The results of the Levenshtein distance test, as seen in Figure 9, suggest that most package hallucinations are not simple off-by-one errors. An off-by-one error in our case refers to a discrepancy between the hallucinated package and its nearest match, involving a difference of 1 to 2 characters, including numbers, letters, or punctuation marks. Our results show that only 13.4% (10,263 of 76,489) have a Levenshtein distance of 1 or 2. Another 37.9% (29,025 of 76,489) of packages registered a score between 3 and 5, which would indicate two words with a common root word or concept that still differ significantly. In particular, 48.6% (37,207 of 76,489) of hallucinations scored 6 or higher, with 20.2% (15,457 of 76,489) of those scoring 10 or higher, indicating two strings that are very different and likely do not share any common theme.
@@
The presence of such a large proportion of high Levenshtein values suggests that the majority of hallucinations are not merely trivial typographical errors but are substantively different from existing package names. The observed results provide further evidence that the root cause of hallucinations is likely to be more complex than minor string manipulation, pointing to deeper issues in the model’s generative processes that govern the creation of package names. The long right tail of the distribution in Figure 9 indicates a wide variety of hallucinations spread over a broad range, revealing a diversity in types of errors and reinforcing that the generation of hallucinations is a complex issue not limited to simple character substitutions, additions, and deletions.
@@
Effect of Deleted Packages. To determine whether packages that existed before a model’s pre-training data cut-off date (i.e., the final day of data included in the model’s training set) but were subsequently removed contribute significantly to package hallucinations, we conducted an analysis using package download counts obtained via Google BigQuery [13]. We searched PyPI download counts from 2022 and earlier to compile a list of packages that existed before 2023 but are no longer available on PyPI. This list was then compared against the master list (obtained from PyPI as of January 10, 2024) of hallucinated packages across all models. We detected 12,871 packages that were available between 2020 and 2022 and have since been removed from PyPI. Of these deleted packages, only 133 (0.17%) were generated during our analysis, indicating that deleted packages are a negligible source of package hallucinations. This finding contradicts our hypothesis, as we expected a sizable percentage of hallucinated packages due to the presence of deleted packages in the training data.
@@
Effect of Language Confusion. Another behavior observed during the main experiment is the tendency to confuse programming languages while generating package output (i.e. the model is asked to provide Python packages, but instead provides JavaScript packages). To validate the potential existence of this behavior in code-generating LLMs, we obtained master lists of packages (using libraries.io [58]) from the nine most popular open-source repositories and compared our list of hallucinated package names generated during Python testing to the respective master lists of valid packages from other programming languages. Any intersection between the two lists indicates a cross-language hallucination. Overall, only JavaScript is a significant source of cross-language hallucinations, as 8.7% (6,705/76,489) of hallucinated Python packages are valid JavaScript packages. All other languages contributed negligible hallucinations, combining for only 0.8% (663 of 76,489) across eight other open-source repositories, including R, Rust, Ruby, PHP, Swift and .NET (see Table 3 for complete results).
@@
RQ4 Summary: Most unique hallucinated package names were generated by a single model and thus only appeared in 1 out of 16 tests. Most hallucinated package names were not semantically similar to a valid Python package as measured by the Levenshtein distance. Deleted packages were a negligible source of package hallucinations, while JavaScript was the only significant source of cross-language hallucinations.
@@ 6 Mitigation
Motivated by the finding of systemic package hallucination in all the tested models, we next investigate techniques to mitigate the occurrence of such hallucinations (thus addressing RQ5).
@@ 6.1 Mitigation Strategies
A straightforward approach to address package hallucinations could be to cross-reference a master list of valid packages with the model’s output, thereby eliminating any incorrect package names. This type of filtering method is ineffective as a defense strategy, as an attacker could immediately publish a hallucinated package to the repository and be subsequently included in the “allow” list. Although using a curated list of “known good packages”, using some metric such as package popularity, would be a more effective method, this is still considered a blunt and reactive approach that requires constant verification and updating. We refer to these types of filtering techniques as post-generation techniques, as they aim to filter hallucinated content after it has been generated by the model. In contrast, the techniques we introduce in this section are pre-generation techniques which aim to prevent the model from generating hallucinated content. These techniques proactively reduce the likelihood of hallucinations at the source with less overhead and maintenance while providing a more reliable output.
Developing specialized mitigation strategies for code-generating LLMs is an area that has not been extensively investigated; therefore, we draw upon general hallucination mitigation strategies designed for typical NLP tasks, which can also be adapted for code generation. These strategies can generally be grouped into two broad categories: prompt engineering and model development [59]. Prompt engineering includes methods such as Retrieval Augmented Generation (RAG), self-refinement, and prompt tuning. RAG approaches involve enriching the original prompt with additional information gathered from an external source, such as the web or a pre-determined database [32]. This augmentation can occur at any stage -- before, during or after response generation -- and can be iterative, improving over multiple cycles until the response is verified to be accurate. Self-refinement strategies, on the other hand, utilize the model itself to detect and refine potential hallucinations.
The second main mitigation strategy involves improving the underlying LLM model itself through improved decoding strategies or supervised fine-tuning. Supervised fine-tuning alters model parameters to improve performance on tasks prone to hallucinations, using a labeled dataset for more precise training. Decoding strategies are also considered a viable mitigation strategy from the literature, but based on our findings from RQ2, we know that altering decoding parameters, such as top-$k$, top-$p$, and min-$p$, do not result in a decreased hallucination rate. We evaluate each of the three remaining categories to determine their applicability to the code generation task and to our specific use case of reducing package hallucinations during code generation.
@@ 6.2 Mitigation Implementation and Results
Retrieval Augmented Generation (RAG). We employ a method of RAG which supplements the prompt with valid package names before generation to assist the model in generating a response. We developed an additional dataset to serve as the supplementary information by taking the top 20,000 most popular PyPI packages and prompting LlaMA-2 to generate a list of five questions that each package could help answer given the description. After removing duplicate responses, this resulted in 65,000 statements in the form “Package [x] could answer questions about [y]”. These 65,000 statements were stored in a vector database, enabling efficient retrieval of semantically similar statements. When a model was asked to recommend packages given a code generation prompt or Stack Overflow question, the vector database is first queried from within the code to return the top 5 most semantically similar statements. These statements are appended to the prompt to give the model additional information containing established valid packages to assist in generating non-hallucinated responses.
Self-Refinement. Drawing on insights from our findings in RQ3 (see Section 5.3), which revealed that LLMs often exhibit proficiency in identifying their own package hallucinations, we implemented a self-refinement method. Following the generation of package names, the model is queried regarding the validity of these packages. If the model indicates that the packages are invalid, the response is regenerated with a specific instruction to not use the invalid package. This regeneration process is allowed to iterate up to five times, acknowledging that many package hallucinations are persistent, as demonstrated in RQ3, and may be generated repeatedly. It is possible during this test that a valid package is misclassified as a hallucination by the model, although because this is an iterative process, the success rate should outweigh the false positive rate over the course of testing, resulting in a net decrease in hallucination rate.
Fine-tuning. For our next method, we fine-tuned the models using the code/package list (Heuristic 1) and prompt/package list (Heuristic 2) pairs that were generated during our initial experiments (Section 5.1). All hallucinations were filtered out and the models were re-trained using the remaining valid responses (560,000 samples). As fine-tuning affects the underlying model weights, we also need to make sure the fine-tuned model retains the ability to produce functional and effective code, which we will test by comparing the code quality of the original models and the fine-tuned ones using the well-known HumanEval benchmark [6]. The HumalEval benchmark is a set of pre-defined prompts and test cases, and the final score is a percentage of problems for which the model generated code is both syntactically correct and passes the test cases, reflecting the model’s ability to produce functional and accurate code.
@@
Results. Overall, all the mitigation strategies we implemented resulted in a reduced rate of package hallucination, with RAG and Supervised Fine-Tuning proving particularly effective (see Table 4). Fine-tuning proved to significantly improve the results, especially for the DeepSeek model, where hallucinations were reduced by 83%, achieving a total rate of just 2.66%, which is a lower rate than any of the ChatGPT models (observed during RQ1). Self-refinement feedback was also much more effective for the DeepSeek model (19% reduction) compared to the CodeLlama model (3% reduction). This aligns with our results in RQ3, where the DeepSeek model was proficient at detecting hallucinations, while the CodeLlama model had a strong bias towards labeling packages as valid, which limited its ability to reliably detect errors. The ensemble method of combining all mitigation strategies further improved the results, reducing hallucination rates by 85% and 64% from their baseline levels for DeepSeek and CodeLlama, respectively.
The code quality of the fine-tuned models, as measured using the HumanEval benchmark tests, did decrease significantly, -26.1% and -3.1% for DeepSeek and CodeLlama respectively, in exchange for substantial improvements in package hallucination rate. Although code quality was negatively affected, fine-tuned scores are still at levels comparable to other high-performing models such as Mistral 7B (26.1%), Llama 65B (23.1%), and Llama 2 7B (12.8%).
In summary, our results demonstrate that while all tested mitigation strategies effectively reduce package hallucinations, fine-tuning comes at the cost of diminished code quality. Further research is needed to develop fine-tuning methods that minimize hallucinations without compromising quality. In the meantime, RAG and self-refinement offer promising alternatives.
@@ 7 Discussion and Conclusion
Since the writing of this paper, several more advanced models have emerged for code generation. These newer models may offer improved performance and different hallucination tendencies and characteristics, which we may have been unable to capture in our study. The study also includes fewer commercial models due to funding constraints, meaning that the findings may not fully represent the performance and hallucination tendencies of the latest commercial LLMs.
In terms of future work, understanding the precise underlying causes of package hallucinations is still an open question. This includes exploring the architecture and components of LLMs that can contribute to these errors, examining the adequacy of tokenizers, and assessing how the composition of the training data and pre-processing impact hallucination tendencies. Identifying and mitigating these underlying issues could lead to more robust and reliable code generation models. Future work could also focus on developing and testing more sophisticated mitigation strategies tailored specifically for code generation tasks. This could involve advanced techniques in prompt engineering, the use of complex knowledge graphs, the refinement of loss functions, and the exploration of new fine-tuning methods. Integrating real-time feedback mechanisms to dynamically adjust model output could further reduce hallucination rates. Understanding how package hallucinations are systemic and persistent at the token level remains crucial.
In conclusion, we systematically studied package hallucinations in code-generating LLMs, including both commercial and open-source models. Our comprehensive analysis revealed that 19.7% of the generated packages are fictitious, posing a critical threat to software security through package confusion attacks. We identified key behavioral patterns and characterized hallucinated packages, proposing effective mitigation strategies. Our findings underscore the importance of addressing package hallucinations to enhance the reliability and security of AI-assisted software development.
@@
We have disclosed our research to model providers including OpenAI, Meta, DeepSeek, and Mistral AI. We received responses from all providers that our research will be taken into consideration for future models.
Our research highlights the feasibility of a new attack vector that can be used to carry out package confusion attacks by exploiting the package hallucinations generated by code-generating LLMs. Hallucinations are a well-studied limitation of generative AI models, including code-generating LLMs, which even current state-of-the-art techniques cannot completely eliminate. Our goal in this research is to get a better understanding of this phenomenon of package hallucinations in code-generating LLMs. We hope that these findings can help secure future models so that users are better protected against package confusion threats enabled due to package hallucinations.
@@
Open Science
We are committed to the open science policy and will make all source code, datasets, and generated code publicly available. The only exception is we will not release our master list of hallucinated package names or the detailed results of our individual tests to the public. Such information could be misused by malicious actors to execute a package hallucination attack. The master list and full results from each of the 30 tests presented in Section 5.1 will be shared responsibly by request to verified researchers.
Specific items that will be made available:
  1. 1.
    All code and data necessary to recreate our primary experiments to detect hallucinated packages.
    • Given a model, a single script will generate all code and packages then evaluate the output for hallucinations and save the results.
    • Both the LLM and Stack Overflow datasets.
    • An environment.yml file that details the packages and dependencies needed to run the code.
  2. 2.
    All code and data to reproduce each figure in the paper.
  3. 3.
    All code and data to reproduce the mitigation tests in the paper.
    • Two fine-tuned models trained on question/answer responses containing valid package names.
    • RAG database along with the code and data used to create it.
    • Code to run all 4 mitigation tests.
  4. 4.
Acknowledgements
This research was partially supported by the University of Texas at San Antonio (UTSA) Faculty Development Leave (FDL) program and by the National Science Foundation (NSF) under award number 2231002.